### v2

@@ Abstract
Motivated by global efforts to develop quantum computing for practical, industrial-scale challenges, we showcase the effectiveness of state-of-the-art hybrid quantum-classical solvers in addressing the business-centric optimization problem of scheduling Automatic Guided Vehicles (AGVs). These solvers leverage a noisy intermediate-scale quantum (NISQ) device, specifically a D-Wave quantum annealer. In our study, the hybrid solvers exhibit non-zero quantum processing times, indicating a significant contribution of the quantum component to solution efficiency. This hybrid methodology performs comparably to existing classical solvers, thus indicating ‘quantum readiness’ for scheduling tasks. Our analysis focuses on a practical, business-oriented scenario: scheduling AGVs within a factory constrained by limited space, simulating a realistic production setting. Our new approach concerns mapping a realistic AGV problem onto a problem reminiscient of railway scheduling and demonstrating that the AGV problem more suits quantum computing than the railway counterpart and is more dense in terms of an average number of constraints per variable. We demonstrate that a scenario involving 15 AGVs, which holds practical significance due to common bottlenecks like shared main lanes leading to frequent deadlocks, can be efficiently addressed by a hybrid quantum-classical solver within seconds. Consequently, our research paves the way for the near-future business adoption of hybrid quantum-classical solutions for AGV scheduling, anticipating that forthcoming improvements in manufacturing efficiency will increase both the number of AGVs deployed and the premium on factory space.
@@
quantum annealing, hybrid quantum-classical optimization, job-shop scheduling, AGV scheduling
@@ 1 Introduction
Quantum computing, a relatively young science field, has developed rapidly over recent years. It is commonly accepted that this novel technology could shape the future. Certain tasks are not tractable with a classical computer; quantum computing approaches are good candidates to handle these. Algorithms such as Shor’s algorithm[1] proposed a polynomial-time quantum algorithm for the integer factoring problem. At its current development level the improved efficiency offered by quantum computing is restricted to certain computational problems [2]. Nevertheless, it is widely recognized that we are in the era of noisy intermediate-scale quantum (NISQ) devices [3, 4]: a few hundred quantum bits are technologically available, but noise and imperfections still have a relevant impact on their operation. Despite that, the first examples of quantum advantage have already been presented[5], and there is an increasing research interest in potential industrial applications that are considered[6, 7, 8]. Currently, the community of quantum computing scientists is searching for practical industrial-scale use cases of industrial problems that can be successfully solved using the available quantum devices [9]. Our intention is to address this quest by tackling the Automated Guided Vehicles (AGV) scheduling problem. The size and characteristics of the scheduling problem addressed in this paper make it a good candidate to explore noisy-intermediate scale quantum (NISQ) [3] technologies in industrial applications.
@@
In the broad picture of production scheduling, an approach considered as promising is Smart Scheduling[11]. It combines the cyber-physical production systems (CPPS) with the decision support system (DSS). The AGV scheduling optimization can be treated either as part of the larger industrial system in light of the philosophy of industry 4.0 [12] (e.g. together with industrial job scheduling) or just as a sole component. For the sake of demonstration, we focus on the latter. However, our model can be re-integrated into a general DSS system.
@@
Many industrial processes can be optimized, including the design of manufacturing systems and processes, assembly line processes, etc. Within these, we address AGV scheduling, in particular, timetabling in a short time horizon. The decisions have to be made almost in real-time, requiring fast computational heuristics. The concept of industry 4.0 [13] implies that AGV scheduling must be tied to the particular factory specifics. To tie our research directly to current business needs, we address a problem that arises in a production environment, i.e. the practice of an actual operating factory. (Its identity and further details are confidential.) We model this particular environment, its configuration, and requirements, addressing the particular needs of the given factory. In this factory, there is a well-defined space where AGVs can maneuver. They are restricted to moving on dedicated "roads" (uni- or bidirectional) to reach ports, loading places, charging stations, etc. Here, the AGVs are controlled by a central system where their scheduling occurs.
@@
AGV scheduling algorithms are divided into rostering111also termed as “scheduling” in some works and routing[14, 15]. The aim of rostering is to dispatch a set of AGVs to perform certain jobs of equipment transportation within the factory. Then, routing (path planning) aims at finding suitable paths for the AGVs, together with the ”timetable” and a possible ordering of AGVs passing certain congested places. These problems are commonly solved via linear programming optimization [15, 16].
@@
The work of Tuan [17] provides a fair overview of the optimization of AGVs scheduling. Vehicle rostering is discussed both as an offline scheduling problem (where all transportation requests are known in advance) or as online scheduling (where environments are stochastic, and requirements appear on the fly, c.f. the work of[18]). We follow the approach of deterministic offline scheduling, but each time the circumstances change, we recalculate schedules. An important constraint (tied to time constraints in scheduling theory) is the maximal window of time in which a task has to be performed by an AGV. The objective is a linear function of variables reflecting completion time costs. To solve both static and dynamic problems, most authors apply ILP together with a (custom) column generation approach and various methods of optimization. As we consider routes and tasks of AGVs predefined for our optimization problem, column generation is not applicable.
The literature on bigger-scale AGV scheduling problems is still scarce, and the considered sizes are rather limited when compared to industry practice. For instance, in a recent paper [19], a system with $9$ AGVs is already considered as a large-scale one. In our use case, we introduce a model tailored to the specific needs of this operational setting, employing proprietary hybrid quantum-classical solvers for resolution. In particular, we concentrate on certain components of AGVs scheduling, AGVs’ timing, and ordering, stressing the problem of deadlock resolution.
@@
When compared to the most similar contributions in the literature, our approach is complementary to the work of Geitz et al. [8], where the job scheduling in a factory is optimized on a quantum annealer or classical device. There, however, the focus is on the factory machines. As opposed to that, we focus on the AGV traffic, assuming job assignments to industrial machines are fixed. Haba et al.[20] also addressed a problem of AGV scheduling with quantum methods. There, however, quantum annealing was applied for the routing of AGVs, as opposed to our problem, which addresses the timing and ordering of the AGVs.
The planning of AGVs’ paths and task assignment, which is widely analyzed along with scheduling in literature [14], is not part of our optimization task. We are required to address AGV scheduling in itself, which has less coverage in the literature. Our contribution is the first one in this research direction, which is dedicated to the particular business-driven problem of AGVs scheduling in a factory with limited space.
@@
This particular optimization problem appears to be complex but still tractable for the meaningful number of AGVs ($15$ AGVs in particular). The anticipated evolution of Industry 4.0, particularly with respect to AGV scheduling, is expected to involve more complex scenarios, such as managing 40 or more AGVs within the confined spaces of a factory. Addressing such challenges in a timely manner could become problematic using common heuristics. Hence, new heuristics or computational paradigms may be necessary for such future applications.
@@
This paper is organized as follows. In Section 3.2, we define our problem and formulate it mathematically. In Section 4, we discuss computational results. In Section 5, we draw conclusions, while Appendix A is devoted to technical details of the mathematical model, and Appendix B is devoted to details on the particular problem of AGVs’ scheduling.
@@ 2 Quantum annealing
A recent theoretical study [21] points toward possible speedup for approximately solving instances of NP-hard combinatorial optimization problems as long as they are implemented using a fault-tolerant quantum computer. Inspired by this, we hope that a quantum-ready model can potentially scale better than classical computer algorithms. Hence, developing quantum hardware may open the possibility of efficiently dealing with large problem instances.
Quantum Annealing [22] is a heuristic algorithm that operates within the principles of Adiabatic Quantum Computing [23, 24], particularly for solving optimization problems. In this regime, a given optimization task is encoded into a physical system described by a problem Hamiltonian $H_{P}$, so that the system’s lowest energy state (ground state) corresponds to the solution of the original problem. Initially, the system is prepared in the ground state of another related physical system, described by the initial Hamiltonian $H_{0}$. Then, the system is slowly evolved into the target system $H_{P}$ by tuning the parameters of the instantaneous Hamiltonian to turn $H_{0}$ into $H_{P}$ in a long enough time $T$. According to the adiabatic theorem [25], if certain conditions hold, the system should remain in the ground state during the evolution and thus reach the ground state of $H_{P}$ at the end, yielding a solution to the original problem. Mathematically, the evolution of the system is described by the following time-dependent Hamiltonian:
@@
In practice, many optimization problems, including our scheduling case, can be modelled as an integer linear program (ILP). The integer variables can be encoded into suitable binaries in various ways [27, 28]. The constraints of the integer programs are taken into account with penalties [29]. The right choice of penalties, a nontrivial problem [30, 31, 32, 33] itself. Although the size of the current quantum annealer is limited in terms of the number of qubits (in our case, approximately $5600$ qubits and $40$ thousands couplings between them organized into the Pegasus topology of QPU [34]), larger optimization problems, that do not fit quantum devices, can be solved via hybrid quantum-classical approach. D-Wave offers two hybrid classical-quantum solvers as a part of its service: BQM and CQM [10, 35].
The BQM solver inputs QUBO problems and solves them with a portfolio of classical heuristics. In the course of the process, certain subproblems are sent to the quantum processor. In the case of the CQM solver, the input is a constrained quadratic program, which can be an ILP. Its transformation to QUBO is performed within the solver. The handling of constraints with penalties is also done by the solver internally.
These solvers are proprietary and closed source; their details are hidden from the users. Nevertheless their operation can be understood according to D-Wave’s white paper [10]. In particular, the data flow is described as follows. The solver reads the input problem. Then, it invokes a portfolio of heuristic solvers that run in parallel using classical CPUs and GPUs in a cloud computing environment. These heuristics contain a Quantum Module (QM), which can send queries to QPU. The quantum results help classical heuristic search to improve the quality of a current pool of solutions. After post-processing (removing duplicates, etc.), the results are forwarded to the user.
@@
3 AGV scheduling
The concept of industry 4.0 [13] implies that AGV scheduling must be tied to the particular factory specifics. To tie our research directly to current business needs, we address a problem that arises in a production environment, i.e. the practice of an actual operating factory. (Its identity and further details are confidential.) We model this particular environment, its configuration, and requirements, addressing the particular needs of the given factory. In this factory, there is a well-defined space where AGVs can maneuver. They are restricted to moving on dedicated "roads" (uni- or bidirectional) to reach ports, loading places, charging stations, etc. Here, the AGVs are controlled by a central system where their scheduling occurs.
Concerning a factory environment with limited space, an important issue with AGVs is the possibility of deadlocks, where several AGVs come to such a point that none of them can move forward because each of them is blocked by the others. [17] mentions various deadlock resolution methods:
@@ 3 AGV scheduling
These deadlock resolution strategies can mostly be encoded as constraints to the ILP. We have opted for the last approach, i.e. dynamic zones. It is worth remembering that deadlock resolution in limited space is challenging and difficult, even with limited problem size (regarding the number of AGVs).
@@
The objective is the weighted sum of completion times [17]. The weights reflect the priorities of AGVs. We assume that AGVs’ tasks and paths (represented by the colour lines in Fig. 1) are pre-defined as input for our problem. Each time these inputs are changed, a new optimization problem is created and recomputed. Such changes in the input can be caused e.g. by disturbances in the AGV traffic.
Figure 1: Example of the AGV scheduling problem given pre-defined AGVs paths as colour lines. First, AGVs are routed to travel between ports, what is the input to our algorithm. The goal of the algorithm is ti timetable AGVs to avoid collisions at zones. In our case zones $s_{0},\ldots s_{6}$ are allegorically assigned spatial areas where AGVs paths intercept start or active ports are located. AGVs paths in terms of zones are AGV $s_{0},s_{1},s_{2},s_{3}$, AGV $s_{4},s_{3},s_{2},s_{1}$, AGV $s_{6},s_{5},s_{4},s_{3}$ $AGVs_{5},s_{6}$ and AGV $s_{2},s_{3}$
@@ 3.1 AGV scheduling: algorithm
Steps $1$ $3$ in processing are pre-processing steps. The higher computational effort is required in step $5$, which deals with optimization. Hence, the analysis of the computation process (both quantum and classical) will refer to step $5$ of the algorithm. Importatnly, the optimization results (values of order variables uniquely defining the order of AGVs at each zone while the traffic beyond zones is conflict-free) can be directly fed into the application programming interface (API) of the operation control software system of the factory we have modeled.
3.2 AGV scheduling: ILP model

### v3

@@ Abstract
Motivated by recent efforts to develop quantum computing for practical, industrial-scale challenges, we demonstrate the effectiveness of state-of-the-art hybrid (not necessarily quantum) solvers in addressing the business-centric optimization problem of scheduling Automatic Guided Vehicles (AGVs). Some solvers can already leverage noisy intermediate-scale quantum (NISQ) devices. In our study, we utilize D-Wave hybrid solvers that implement classical heuristics with potential assistance from a quantum processing unit. This hybrid methodology performs comparably to existing classical solvers. However, due to the proprietary nature of the software, the precise contribution of quantum computation remains unclear.
Our analysis focuses on a practical, business-oriented scenario: scheduling AGVs within a factory constrained by limited space, simulating a realistic production setting. Our approach maps a realistic AGVs problem onto one reminiscent of railway scheduling and demonstrates that the AGVs problem is better suited to quantum computing than its railway counterpart, the latter being denser in terms of the average number of constraints per variable. The main idea here is to highlight the potential usefulness of a hybrid approach for handling AGVs scheduling problems of practical sizes. We show that a scenario involving up to $21$ AGVs, significant due to possible deadlocks, can be efficiently addressed by a hybrid solver in seconds.
@@
quantum annealing, hybrid quantum-classical optimization, job-shop scheduling, AGVs scheduling
@@ 1 Introduction
Quantum computing, a relatively young science field, has developed rapidly in recent years. It is commonly accepted that this novel technology could shape the future. Certain tasks are not tractable with a classical computer; quantum computing approaches are good candidates to handle these. For example, Shor’s algorithm[1] proposes a polynomial-time quantum algorithm for the integer factoring problem. At its current development level, the improved efficiency offered by quantum computing is restricted to certain computational problems [2]. Nevertheless, it is widely recognized that we are in the era of noisy intermediate-scale quantum (NISQ) devices [3, 4]: a few hundred qubits are technologically available, but noise and imperfections still have a relevant impact on their operation. Despite that, the first examples of quantum computing utility have already been presented[5] for some specific physical systems, and there is an increasing research interest in potential industrial applications that are considered[6, 7, 8]. Currently, the community of quantum computing scientists is searching for practical industrial-scale use cases of industrial problems that can be successfully solved using the available quantum devices [9]. Our intention is to address this quest by tackling the Automated Guided Vehicles (AGVs) scheduling problem. The size and characteristics of the scheduling problem addressed in this paper make it a good candidate to explore noisy-intermediate scale quantum (NISQ) [3] technologies in industrial applications.
@@
In the broad picture of production scheduling, an approach considered as promising is Smart Scheduling[11]. It combines the cyber-physical production systems (CPPS) with the decision support system (DSS). The AGVs scheduling optimization can be treated either as part of the larger industrial system in light of the philosophy of industry 4.0 [12] (e.g. together with industrial job scheduling) or just as a sole component. For the sake of demonstration, we focus on the latter. However, our model can be re-integrated into a general DSS system.
@@
Many industrial processes can be optimized, including the design of manufacturing systems and processes, assembly line processes, etc. Within these, we address AGVs scheduling, in particular, timetabling in a short time horizon. The decisions have to be made almost in real-time, requiring fast computational heuristics. The concept of industry 4.0 [13] implies that AGVs scheduling must be tied to particular factory characteristics. To tie our research directly to current business needs, we address a problem that arises in a production environment, i.e. the practice of an actual operating factory. (Its identity and further details are confidential.) We model this particular environment, its configuration, and requirements, addressing the particular needs of the given factory. In this factory, there is a well-defined space where AGVs can maneuver. They are restricted to moving on dedicated "roads" (uni- or bidirectional) to reach ports, loading places, charging stations, etc. Here, the AGVs are controlled by a central system where their scheduling occurs.
@@
AGVs scheduling algorithms are divided into rostering 111also termed as “scheduling” in some works and routing[14, 15]. The aim of rostering is to dispatch a set of AGVs to perform certain jobs of equipment transportation within the factory. Then, routing (path planning) aims at finding suitable paths for the AGVs, together with the ”timetable” and a possible ordering of AGVs passing certain congested places. These problems are commonly solved via linear programming optimization [15, 16].
@@
The work of Tuan [17] provides a fair overview of the optimization of AGVs scheduling. Vehicle rostering is discussed both as an offline scheduling problem (where all transportation requests are known in advance) or as online scheduling (where environments are stochastic, and requirements appear on the fly, c.f. the work of[18]). We follow the approach of deterministic offline scheduling, but each time the circumstances change, we recalculate schedules. An important constraint (tied to time constraints in scheduling theory) is the maximal window of time in which an AGV has to perform a task. The objective is a linear function of variables reflecting completion time costs. To solve both static and dynamic problems, most authors apply ILP together with a (custom) column generation approach and various methods of optimization. As we consider routes and tasks of AGVs predefined for our optimization problem, column generation is not applicable.
The literature on bigger-scale AGVs scheduling problems has been developed in recent years to catch up with the increasing size of problems in industry practice. For instance, in a paper from 2019 [19], a system with $9$ AGVs had been considered as a large-scale one, while in a more recent paper [20] from 2023 classical heuristics for much larger AGVs problem were introduced. (Therein, the largest problem was of a few hundred AGVs, but with a computational time of days. The problem solved in a reasonable time of seconds concerns 30-40 AGVs with a particular problem layout.) In our use case, we introduce a model tailored to the specific needs of the real-life-driven operational setting, employing proprietary hybrid quantum-classical solvers for AGVs problems of sizes similar to the above-mentioned. In particular, we concentrate on certain components of AGVs scheduling, AGVs timing , and ordering, stressing the problem of deadlock resolution. As for the actual problem size that a quantum approach can handle, scheduling 2 to 8 AGVs for 10 - 11 tasks (together with routing) on an optical coherent Ising machine was recently reported by Tang et al.[21] (during the preparation of our present contribution). The experimental results therein show that the optical quantum computer has an advantage in computational time over the traditional state-of-art approach.
@@
When compared to the most similar contributions in the literature, our approach is complementary to the work of Geitz et al. [8], where the job scheduling in a factory is optimized on a quantum annealer or classical device. There, however, the focus is on the factory machines. In contrast, we focus on the AGVs traffic, assuming job assignments to industrial machines are fixed. Haba et al.[22] also addressed a problem of AGVs scheduling with quantum methods. There, however, quantum annealing was applied to the routing of AGVs, as opposed to our problem, which addresses the timing and ordering of the AGVs.
The planning of AGVs paths and task assignment, which is widely analyzed along with scheduling in literature [14], is not part of our optimization task. We are required to address AGVs scheduling in itself, which has less coverage in the literature. Our contribution is the first one in this research direction, which is dedicated to the particular business-driven problem of AGVs scheduling in a factory with limited space. In detail, we introduce a new model for AGVs scheduling (ordering), which is simple enough to be handled by a hybrid quantum-classical approach in a competitive manner and is complex enough to have practical significance, as it is derived from a practical use case. Our high-level goal is to pave the way for the development of new open-source hybrid solvers by the application of the most promising, though proprietary, closed-source D-Wave hybrid solver. There have been proposals to develop hybrid solvers for scheduling optimization problems[23, 24]. In these contributions, the optimization task is formulated as a Quadratic Unconstrained Binary Optimization (QUBO) problem, and then its relaxed versions or some subproblems are solved using a quantum annealer, while the remaining part is solved using classical heuristics. However, the transformation of a linear problem into the quadratic representation of QUBO creates a large overhead and requires tuning QUBO parameters. In the present contribution, we use a hybrid solver that inputs the native linear programming model and changes only smaller sub-problems into the QUBO if necessary, which we expect to be more efficient for problems that have suitable linear models, such as AGVs scheduling.
@@
The particular optimization problem handled in our paper appears to be complex but it still tractable for the meaningful number of AGVs ($21$ AGVs in particular). The anticipated evolution of Industry 4.0, particularly with respect to AGVs scheduling, is expected to involve more complex scenarios, such as managing 40 or more AGVs within the confined spaces of a factory. Addressing such challenges in a timely manner could become problematic using common heuristics. Hence, new heuristics or computational paradigms may be necessary for such future applications.
@@
This paper is organized as follows. In Section 2, we discuss Quantum Annealing. In Section 3, we define our problem and formulate it mathematically. In Section 4, we discuss computational results. In Section 5, we draw conclusions, while Appendix A is devoted to details on the particular problem of AGVs scheduling.
@@ 2 Quantum annealing
Quantum Annealing [25] is a heuristic algorithm that operates within the principles of Adiabatic Quantum Computing [26, 27], particularly for solving optimization problems. In this regime, a given optimization task is encoded into a physical system described by a problem Hamiltonian $H_{P}$, so that the system’s lowest energy state (ground state) corresponds to the solution of the original problem. Initially, the system is prepared in the ground state of another related physical system, described by the initial Hamiltonian $H_{0}$. Then, the system slowly evolved into the target system $H_{P}$ by tuning the parameters of the instantaneous Hamiltonian to turn $H_{0}$ into $H_{P}$ in a long enough time $T$. According to the adiabatic theorem [28], if certain conditions hold, the system should remain in the ground state during the evolution and thus reach the ground state of $H_{P}$ at the end, yielding a solution to the original problem. Mathematically, the evolution of the system is described by the following time-dependent Hamiltonian:
@@
In practice, many optimization problems, including our scheduling case, can be modelled as an integer linear program (ILP). The integer variables can be encoded into suitable binaries in various ways [30, 31]. The constraints of the integer programs are taken into account with penalties [32]. The right choice of penalties is a nontrivial problem [33, 34, 35, 36] itself. Although the size of the current quantum annealer is limited in terms of the number of qubits (in our case, approximately $5600$ qubits and $40$ thousands couplings between them organized into the Pegasus topology of QPU [37]), larger optimization problems, that do not fit quantum devices, can be solved via hybrid quantum-classical approach. D-Wave offers two hybrid classical-quantum solvers as a part of its service: BQM and CQM [10, 38].
The BQM solver inputs QUBO problems and solves them with a portfolio of classical heuristics. In the course of the process, certain subproblems are sent to the quantum processor. In the case of the CQM solver, the input is a constrained quadratic program, which can be an ILP. Its transformation to QUBO is performed within the solver. The solver also handles constraints with penalties internally.
These solvers are proprietary and closed source; their details are hidden from the users. Nevertheless, their operation can be understood according to D-Wave’s white paper [10]. In particular, the data flow is described as follows. The solver reads the input problem. Then, it invokes a portfolio of heuristic solvers that run in parallel using classical CPUs and GPUs in a cloud computing environment. These heuristics contain a Quantum Module (QM), which can send queries to QPU. The quantum results help classical heuristic search to improve the quality of a current pool of solutions. After post-processing (removing duplicates, etc.), the results are forwarded to the user.
@@
3 AGVs scheduling
The concept of industry 4.0 [13] implies that AGVs scheduling must be tied to the particular factory specifics. To tie our research directly to current business needs, we address a problem that arises in a production environment, i.e. the practice of an actual operating factory. (Its identity and further details are confidential.) We model this particular environment, its configuration, and requirements, addressing the particular needs of the given factory. In this factory, there is a well-defined space where AGVs can maneuver. They are restricted to moving on dedicated "roads" (uni- or bidirectional) to reach ports, loading places, charging stations, etc. Here, the AGVs are controlled by a central system where their scheduling occurs.
Concerning a factory environment with limited space, an important issue with AGVs is the possibility of deadlocks, where several AGVs come to such a point that none of them can move forward because the others block each of them. [17] mentions various deadlock resolution methods:
@@ 3 AGVs scheduling
These deadlock resolution strategies can mostly be encoded as constraints to the ILP. We have opted for the last approach, i.e. dynamic zones. It is worth remembering that deadlock resolution in limited space is challenging, even with a limited problem size (regarding the number of AGVs).
@@
The objective is the weighted sum of completion times [17]. The weights reflect the priorities of AGVs. We assume that AGVs’ tasks and paths (represented by the colour lines in Fig. 1) are pre-defined as input for our problem. Each time these inputs are changed, a new optimization problem is created and recomputed. Such changes in the input can be caused, for example, by disturbances in the AGVs’ traffic.
Figure 1: Example of the AGVs scheduling problem given pre-defined AGVs paths as colour lines. First, AGVs are routed to travel between ports (colour rectangles), which is the input to our algorithm. The goal of the algorithm is to timetable AGVs to avoid collisions in zones. In our case zones $s_{0},\ldots s_{6}$ are allegorically assigned spatial areas where AGVs paths intercept start or active ports are located. AGVs paths in terms of zones are AGV $s_{0},s_{1},s_{2},s_{3}$, AGV $s_{4},s_{3},s_{2},s_{1}$, AGV $s_{6},s_{5},s_{4},s_{3}$ $AGVs_{5},s_{6}$ and AGV $s_{2},s_{3}$
@@ 3.1 AGVs scheduling: algorithm
Steps $1$ $3$ in processing are pre-processing steps. The higher computational effort is required in step $5$, which deals with optimization. Hence, the analysis of the computation process (both quantum and classical) will refer to step $5$ of the algorithm. Importantly, the optimization results (values of order variables uniquely defining the order of AGVs at each zone while the traffic beyond zones is conflict-free) can be directly fed into the application programming interface (API) of the operation control software system of the factory we have modelled.
3.2 AGVs scheduling: ILP model